How does the shape of earth affect the value of “g” ?

Grade:12th Pass

2 Answers

Shreya gupta
6 years ago
An earth has a geoid shape.It is flattenedAt the poles nd bulges out at the equator.The equitorial radius is 21 km greater than at the pole.g=GM/R^2gravity of earth(ge) = GMe/(Radius of earth )^2ge is inversely proportional to the square of the radius of the earthAs the radius increases the g decreasesR equitorial > R poleg pole > g equitorial
Arun
6 years ago
as it is considered that the earths shape is oval, the distance from the centre is varied.As we know that the gravity is the pull by the earth`s centre, it varies as the distance from the centre varies and thus the shape of earth makes gravity vary.